OpenStudy (anonymous):

ln(y+6) - ln(4)= x+ln(x); solve for y... please help??(:

OpenStudy (ankit042):

write x + ln(x) as lne^x +lnx Now use lna+lnb = ln(a*b) and lna-lnb=ln(a/b)

OpenStudy (anonymous):

so, ln (y+6)/4 = ln(x)+ ln(x) ..?

OpenStudy (ankit042):

NO ln (y+6)/4 = ln (e^x*x)